OpenStudy (anonymous):

The function f(x) = 15(2)x represents the growth of a frog population every year in a remote swamp. Elizabeth wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Elizabeth's purposes? f(x) = 15(2 to the one half power)2x f(x) = 15(22)the x over 2 power f(x) = fifteen halves(2)x f(x) = 30(2)x

OpenStudy (anonymous):

@Gabylovesyou lmao same name

OpenStudy (anonymous):

15(_2^1/2)^2x 15(2^2)^x/2 15/2(2)^x 30(2)^x

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

15(2)^x=15(2,1/2)^2x

OpenStudy (anonymous):

Why? @$w3G_Godd

OpenStudy (anonymous):

Do you think you can help me on another problem?

OpenStudy (anonymous):

ya ill try

OpenStudy (anonymous):

Izzy's heartbeat rate was measured on an EKG machine. After conversion, the function produced could be modeled by a sine function, and the wave produced a maximum of 8, minimum of −2, and period of ð. Which of the following functions could represent Izzy's EKG read-out? f(x) = 8 sin ðx − 2 f(x) = 5 sin ðx + 3 f(x) = 5 sin 2x + 3 f(x) = 2 sin 2x + ð

OpenStudy (anonymous):

maximum of the A function is 8*1-2 = 6, so it can't be. maximum of the D function is 2*1 +pi ~ 5, so it can't be either. Both b and c are viable. By definition of periodic function, if f(x) has period T, f(kx) has period T/k, it means that the period of function b would be 2, and function c would be pi, which is the required answer.

OpenStudy (anonymous):

So C would be the correct answer?

OpenStudy (anonymous):

@$w3G_Godd